The television show Pretty Betty has been successful for many years. That show recently had a share of 15, meaning that among the TV sets in use, 15% were tuned to Pretty Betty. Assume that an advertiser wants to verify that 15% share value by conducting its own survey, and a pilot survey begins with 11 households have TV sets in use at the time of a Pretty Betty broadcast.

Find the probability that none of the households are tuned to Pretty Betty.
P(none) =

Find the probability that at least one household is tuned to Pretty Betty.
P(at least one) =

Find the probability that at most one household is tuned to Pretty Betty.
P(at most one) =

If at most one household is tuned to Pretty Betty, does it appear that the 15% share value is wrong? (Hint: Is the occurrence of at most one household tuned to Pretty Betty unusual?)

  • no, it is not wrong
  • yes, it is wrong
Answer #1

a)
Here, n = 11, p = 0.15, (1 - p) = 0.85 and x = 0
As per binomial distribution formula P(X = x) = nCx * p^x * (1 - p)^(n - x)

We need to calculate P(X = 0)
P(X = 0) = 11C0 * 0.15^0 * 0.85^11
P(X = 0) = 0.1673

b)
P(at least one) = 1 - 0.1673 = 0.8327

c)
P(X <= 1) = (11C0 * 0.15^0 * 0.85^11) + (11C1 * 0.15^1 * 0.85^10)
P(X <= 1) = 0.1673 + 0.3248
P(X <= 1) = 0.4921

d)
no, it is not wrong